OpenStudy (anonymous):

√28 + 4√63 - 2√7 simplify assume that all the variables are positive

OpenStudy (johnweldon1993):

Hint \[\large \sqrt{28} = \sqrt{7 \times 4}\] \[\large \sqrt{63} = \sqrt{9 \times 7}\]